# U-Motion: Learned Point Cloud Video Compression with U-Structured Motion Estimation

### v1

@@ Abstract
Point cloud video (PCV) is a versatile 3D representation of dynamic scenes with many emerging applications. This paper introduces U-Motion, a learning-based compression scheme for both PCV geometry and attributes. We propose a U-Structured multiscale inter-frame prediction framework, U-Inter, which performs layer-wise explicit motion estimation and compensation (ME/MC) at different scales with varying levels of detail. It integrates both higher and lower-scale motion features, in addition to the information of current and previous frames, to enable accurate motion estimation at the current scale. In addition, we design a cascaded spatial predictive coding module to capture the inter-scale spatial redundancy remaining after U-Inter prediction. We further propose an effective context detach and restore scheme to reduce spatial-temporal redundancy in the motion and latent bit-streams and improve compression performance. We conduct experiments following the MPEG Common Test Condition and demonstrate that U-Motion can achieve significant gains over MPEG G-PCC-GesTM v3.0 and recently published learning-based methods for both geometry and attribute compression.
@@ 1 Introduction
One essential element in point cloud video compression is inter-frame coding, which exploits the temporal redundancy between consecutive frames. This is typically accomplished through motion estimation and compensation. However, compared with 2D videos on a regular 2D grid, the points in a PCV are irregularly spread over a small fraction of the 3D space, making the design of motion estimation and compensation approaches for PCV significantly more challenging. Moreover, to ensure superior rate-distortion performance, we need to incorporate efficient context-based coding schemes to compress the estimated motion fields as well as features for recovering the geometry or attributes. Furthermore, it is desirable to develop a codec framework that can be used for both the geometry and attributes, to save hardware design resources for future chip tape-outs. .
@@
To tackle the challenge of motion estimation in point cloud videos, we propose a novel unified point cloud compression scheme, named U-Motion, which incorporates an U-Structured inter prediction module, named U-Inter, to estimate and compensate motion at different scales. The U-Inter structure, akin to the popular U-Net, through the use of both top-down, bottom-up and skip connections, leverages the motion information at both upper and lower scales to inform motion estimation at a current scale. We also propose novel network structures for context detach and context restore for compressing motion and frame features, using the coarser motion as the context for coding the current motion, and using both inter-context (motion-compensated from the reference frame) and intra-context (upsampled from lower scale) for coding the frame features.
We evaluate our method following the MPEG Common Test Conditions (CTC)’s specification [1] for data splitting, in both geometry and color coding. The experimental results show that U-Motion outperforms the learning-based state-of-the-art Unicorn [32], as well as the rule-based MPEG standard G-PCC-GesTM-v3.0 [6] by a large margin in terms of rate-distortion (R-D) trade-off. We conduct ablation studies to validate the effectiveness of each component in our proposed U-Motion codec, and show that the proposed method is robust to different types of human point clouds and motion patterns. Our contributions can be summarized as follows:
  • We propose a novel unified PCV compression scheme, named U-Motion, which incorporates a hierarchical motion estimation module using a U-Net like structure. We demonstrate the importance of both top-down and bottom-up motion calculations as well as skip connections for accurate motion estimation and efficient motion coding under rate constraint. The framework can be used for both geometry and attribute coding.
  • We compress the motion feature by contextually detaching the estimated motion at a lower layer, and conduct a 2-stage temporal-and-spatial context detach to compress the attribute feature.
  • We evaluate our proposed U-Motion codec following the MPEG CTC’s data splitting, for both geometry coding and color coding, and demonstrate U-Motion’s significant performance gain against the learning-based state-of-the-art method Unicorn [32] and MPEG standard G-PCC-GesTM v3.0 [6] in terms of rate-distortion trade-off.
@@ 2.1 Traditional Point Cloud Video Compression
Figure 1: The overall architecture of U-Motion for color attribute coding. The dashed line denotes context connection for entropy coding. Detailed architecture of each module is provided in supplementary material.
@@ 2.2 Learning-Based Point Cloud Video Compression
Early research on learning-based point cloud compression focuses on static point clouds [28, 29, 10, 34, 11]. Wang et al. [28] first propose to use sparse convolution [5] to compress voxelized point clouds, and propose a multiscale architecture, as well as a local auto-regressive architecture SOPA [29] for lossless compression. To further exploit the temporal redundancy in point cloud video, some methods adopt temporal convolution for inter prediction. Anique et al. [2] design a predictor for geometry compression, which compensates and compresses the current frame’s latent representation based on its reference neighbors. Wang et al. introduce Unicorn [31, 32], a unified architecture that integrates geometry & attributes, static & dynamic compression, using context compensation and multiscale motion compensation. The temporal compensation is obtained through sparse convolution on the voxels in the previous frame, known as target convolution. However, both Unicorn and Anique et al.’s method lack explicit motion estimation, leading to sub-optimal temporal prediction. On the contrary, Fan et al. propose D-DPCC [9] with an end-to-end learning-based motion estimation and motion compensation (ME/MC) module. Xia et al. [33] further propose a two-layer ME/MC architecture for coarse-to-fine motion compensation. However, the above ME/MC-based methods are developed only for point cloud video geometry. In addition, their lack of a hierarchical inter- and intra-prediction structure leads to inferior coding efficiency compared to Unicorn-Geometry.
2.3 Leaning-based Video Compression
Prior to learning-based point cloud compression, there have been significant advances in learned video compression [22, 15, 16, 18, 26, 19, 20, 21]. DVC [22] and FVC [15] are the pioneering AI video codecs that performs ME/MC on pixel domain and feature domain, respectively. These work follow a residual-coding framework which encodes the difference between the warped decoded reference frame and the current frame. Li et al. propose DCVC [18], which extends the residual-coding framework to context-coding, generating temporal context priors instead of warped predictions for ME/MC. DCVC-HEM [19] further proposes a parallel checkerboard entropy model supporting variable bit-rate. DCVC-DC [20] and DCVC-FM [21] further incorporate hierarchical quality structure and inter feature modulation to alleviate the accumulation of error in the temporal context. However, due to the irregular point geometry and data diversity of point cloud sequences, learning-based video compression methods cannot be directly integrated to point cloud video compression.
@@ 3.1 Overview of Attribute Compression
As shown in Figure 1, our method is a hierarchical architecture consisting of $(H-L+1)$ U-Inter and Spatial Predictive Coding Layers, where the $l$-th layer corresponds to the $l$-th level of the point cloud when represented by an octree. We denote the point cloud frame (including both geometry coordinate and color attribute for each point at each octree level) at time $t$ and layer $l$ as $p_{l}^{t}$, and its reference (i.e. the decoded previous point cloud frame) as $\hat{p}_{l}^{t-1}$. $p_{l}^{t}$ is expressed as a sparse tensor $\left[G(p_{l}^{t}),A(p_{l}^{t})\right]$, where $G(p_{l}^{t})$ denotes an $N\times 3$ geometry coordinate matrix and $A(p_{l}^{t})$ denotes the corresponding $N\times 3$ attribute matrix, with $N$ denoting the total number of points at layer $l$. The encoder fuses $p_{l}^{t}$ and the upper layer’s latent $f_{l+1}^{t}$ to generate the latent representation $f_{l}^{t}$. The same encoder is also used to generate the reference frame’s latent $\hat{f}_{l}^{t-1}$.
The U-Net structured U-Inter module applies both top-down and bottom-up inter prediction strategies. The top-down motion feature extraction infers the initial motion embedding $e_{d,l}^{t}$ between $f_{l}^{t}$ and $f_{l}^{t-1}$, by exploiting both $f_{l}^{t}$ and $\hat{f}_{l}^{t-1}$, as well as the higher-scale motion information $e_{d,l+1}^{t}$. The bottom-up inter prediction further contextually detach $e_{u,l-1}^{t}$ from $e_{d,l}^{t}$ to generate motion embedding $e_{l}^{t}$. Essentially, it removes the redundancy of $e_{d,l}^{t}$ conditioned on $e_{u,l-1}^{t}$. The motion compression module compresses and decompresses $e_{l}^{t}$ into $\hat{e}_{l}^{t}$, which is then contextually restored (Section 3.4) into $e_{u,l}^{t}$ with $e_{u,l-1}^{t}$ as context. The motion decoding module decodes the motion field $m_{l}^{t}$ from $e_{u,l}^{t}$ and the lower layer motion $m_{l-1}^{t}$ Finally, the network performs motion compensation to generate inter context $\check{f}_{l}^{t}$ with $m_{l}^{t}$ by warping and interpolation.
The Spatial Predictive Coding module further reduces the layer-wise spatial redundancy by generating an intra context $\bar{f}_{l}^{t}$ based on the already decoded lower layer’s latent $\hat{f}_{l-1}^{t}$. The inter context $\check{f}_{l}^{t}$ and intra context $\bar{f}_{l}^{t}$ are then contextually detached from $f_{l}^{t}$ to generate residual $r_{l}^{t}$. The latent compression module compresses and decompresses $r_{l}^{t}$ into $\hat{r}_{l}^{t}$, which is then combined with the context $\bar{f}_{l}^{t}$ and $\check{f}_{l}^{t}$ to contextually restore the decoded latent $\hat{f}_{l}^{t}$. The reconstruction module takes $\hat{f}_{l}^{t}$ and the previous layer’s latent $\hat{p}_{l-1}^{t}$ to recover the color attributes in $\hat{p}_{l}^{t}$.
3.2 Hierarchical Attribute Encoding and and Reconstruction
@@
The encoder maps the current layer $p_{l}^{t}$ and the latent from the upper layer $f_{l+1}^{t}$ into the latent representation for the current layer $f_{l}^{t}$. The same encoder is used to map $\hat{p}_{l}^{t-1}$ and $\hat{f}_{l+1}^{t-1}$ into $\hat{f}_{l}^{t-1}$. In contrast to Unicorn which blocks the latent from flowing between layers, our encoder fuses the higher layer’s latent $f_{l+1}^{t}$, in the hope that the lower layer’s latent representation encapsulate important information of the higher layers. The convolution indicated in this module refers to sparse convolution, as with all other modules of the entire network.
Attribute Reconstruction.
The reconstruction module recovers the colors in the $l$-th scale point cloud $\hat{p}_{l}^{t}$ based on the reconstructed latent $\hat{f}_{l}^{t}$ and lower scale’s reconstruction $\hat{p}_{l-1}^{t}$. For attribute compression with geometry coordinates known, the network estimates a residual attribute $\Delta p_{l}^{t}$ every layer based on the concatenation of upsampled $\hat{p}_{l-1}^{t}$ and $\hat{f}_{l}^{t}$. The reconstruction is the sum of $\Delta p_{l}^{t}$ and the upsampled $\hat{p}_{l-1}^{t}$. The upsample of color is realized by trans-pooling.
3.3 U-Inter for ME/MC
@@
As shown in Figure 1, our U-Inter Prediction follows a U-Net structure, consisting of top-down and bottom-up connections between scales. At an intermediate scale (e.g. $l=L+1$ in the figure), the motion field $m_{l}^{t}$ is generated using both the current latent $f_{l}^{t}$ and reference latent $\hat{f}_{l}^{t}$, as well as the upper-scale motion embedding $e_{d,l+1}^{t}$, and the lower-scale motion embedding $e_{u,l-1}^{t}$ and motion flow $m_{l-1}^{t}$.
Figure 2: The network architecture for context detach and restore.
Specifically, the motion feature extraction module takes the concatenation [9] of $f_{l}^{t}$ and $\hat{f}_{l}^{t}$, as well as the upper-scale flow embedding $e_{d,l+1}^{t}$ (down-sampled to the current scale using a stride 2 convolution) as input and generates the initial motion embedding $e_{d,l}^{t}$ of the current layer. The top-down connection to $e_{d,l+1}^{t}$ enables $e_{d,l}^{t}$ to inherit the motion features extracted from the upper scale. However, compressing $e_{d,l}^{t}$ directly is suboptimal because 1) $e_{d,l}^{t}$ may be similar across layers; and 2) $e_{d,l}^{t}$ has a limited receptive field due to the lack of lower-scale information. Therefore, the bottom-up inter prediction contextually detaches the redundancy in $e_{d,l}^{t}$ that is predictable from the already decoded lower-scale $e_{u,l-1}^{t}$. The output after detaching, i.e. the motion embedding $e_{l}^{t}$ is compressed and decompressed into $\hat{e}_{l}^{t}$. $\hat{e}_{l}^{t}$ is then contextually restored into $e_{u,l}^{t}$ and fused with lower-scale motion information using lower-scale motion $e_{u,l-1}^{t}$ as context. Finally, $e_{u,l}^{t}$ and the estimated motion from the lower layer $m_{l-1}^{t}$ (after upsampling) are input to the motion decoder to generate the estimated motion $m_{l}^{t}$.
Prior approaches for motion estimation often adopt a coarse-to-fine approach, starting with estimating the motion from the lowest scale, and then progressively determining the residual motion in each upper scale. We have observed, though, the lower scale point clouds may lack details that are necessary for estimating the motion correctly. When the top-down links in the U-Inter structure were removed, the network often yield completely zero motions, in part due to the rate constraint. The top-down connection to $e_{d,l+1}^{t}$ and the bottom-up connection to $e_{u,l-1}^{t}$ and $m_{l-1}^{t}$, through multiple scales, enable the proposed U-Inter architecture to estimate large motion effectively, by inheriting the details of upper layers for lower layer motion estimation.
Figure 3: The architecture of Spatial Predictive Coding layer for Geometry. The dashed line denotes context connection. The long dashed line denotes the action of assigning the decoded geometry coordinate $\hat{p}_{l}^{t}$ to U-Inter to get $\check{f}_{l}^{t}$.
Motion Decoding and Compensation.
We extend the single-scale motion compensation in prior works to multi-scale to leverage the details at higher layers. To get the $l$-th scale motion for compensation, the motion decoding module decodes a residual motion flow $\Delta m_{l}^{t}$ from both motion feature $e_{u,l}^{t}$ and upsampled $m_{l-1}^{t}$, which is then added to the upsampled $m_{l-1}^{t}$ to generate the fine motion $m_{l}^{t}$. For each point with coordinate $s$ in $f_{l}^{t}$, its warped position $\check{s}=s+m_{l,s}^{t}$ is likely to be invalid in $\hat{f}_{l}^{t-1}$. Therefore, we adopt the 3D Adaptive Weighted Interpolation (3DAWI) algorithm [9] to generate the inter context $\check{f}_{l}^{t-1}$,
$\check{f}_{l,s}^{t}=\frac{\sum_{v\in\vartheta(\check{s})}d_{v}^{-1}\cdot\hat{f}_{l,v}^{t-1}}{\max\left\{\sum_{v\in\vartheta(\check{s})}d_{v}^{-1},\alpha\right\}},$ (1)
@@
where $\vartheta(\check{s})$ denotes the K-nearest neighbors of $\check{s}$ in $\hat{f}_{l}^{t-1}$, $d_{v}$ is the distance between $\check{s}$ and $v$, $\alpha$ is a penalty parameter to decrease the weights of the points too far from $\check{s}$. Different from 3DAWI, We make $\alpha$ a learnable parameter for more flexibility. During experiments, we set $K=6$ and initial $\alpha$ to $4$. $\alpha$ converges to $3.6$ when training stops.
3.4 Spatial Predictive Coding

### v2

# U-Motion: Learned Point Cloud Video Compression with U-Structured Temporal Context Generation

@@ Abstract
Point cloud video (PCV) is a versatile 3D representation of dynamic scenes with emerging applications. This paper introduces U-Motion, a learning-based compression scheme for both PCV geometry and attributes. We propose a U-Structured inter-frame prediction framework, U-Inter, which performs explicit motion estimation and compensation (ME/MC) at different scales with varying levels of detail. It integrates Top-Down (Fine-to-Coarse) Motion Propagation, Bottom-Up Motion Predictive Coding and Multi-scale Group Motion Compensation to enable accurate motion estimation and efficient motion compression at each scale. In addition, we design a multi-scale spatial-temporal predictive coding module to capture the cross-scale spatial redundancy remaining after U-Inter prediction. We conduct experiments following the MPEG Common Test Condition for dense dynamic point clouds and demonstrate that U-Motion can achieve significant gains over MPEG G-PCC-GesTM v3.0 and recently published learning-based methods for both geometry and attribute compression.
@@ 1 Introduction
One essential element in point cloud video compression is inter-frame prediction, which exploits the temporal redundancy between consecutive frames. This is typically accomplished through motion estimation and compensation. However, compared to 2D videos on a regular 2D grid, the points in a PCV are irregularly spread over a small fraction of the 3D space, making the design of motion estimation and compensation approaches for PCV significantly more challenging. Moreover, to ensure superior rate-distortion performance, we need to incorporate efficient context-based coding schemes to compress the estimated motion fields as well as features for recovering the geometry or attributes. Furthermore, it is desirable to develop a codec framework that can be used for both the geometry and attributes, to save hardware design resources for future chip tape-outs. .
@@
To tackle the challenge of motion estimation in point cloud videos, we propose a novel unified point cloud compression scheme, named U-Motion, which incorporates an U-Structured inter prediction module, named U-Inter, to estimate and compensate motion at different scales. The U-Inter structure, akin to the popular U-Net, through the use of both top-down motion propagation, bottom-up motion predictive coding and skip connections, leverages the motion information at both upper and lower scales to inform motion estimation and coding at a current scale. U-Inter also incorporates a multi-scale group motion compensation scheme, utilizing the hierarchically reconstructed motion flow to enrich temporal context generation.
We evaluate our method following the MPEG Common Test Conditions (CTC), dense dynamic part’s specification [1] for data splitting, in both geometry and color coding. The experimental results show that U-Motion outperforms the learning-based state-of-the-art Unicorn [33], as well as the rule-based MPEG standard G-PCC-GesTM-v3.0 [6] by a large margin in terms of rate-distortion (R-D) trade-off. We conduct ablation studies to validate the effectiveness of each component in our proposed U-Motion codec. Our contributions can be summarized as follows:
  • We propose a novel unified PCV compression scheme, named U-Motion, which incorporates a hierarchical U-Inter Prediction module using a U-Net like structure. The proposed U-Inter incorporates Top-Down Motion Propagation, Bottom-Up Motion Predictive Coding and Multi-scale Group Motion Compensation for accurate motion generation and efficient motion coding under rate constraint. The framework can be used for both geometry and attribute coding.
  • We evaluate the U-Motion codec following the MPEG CTC’s data splitting, for both geometry coding and color coding, and demonstrate U-Motion’s significant performance gain against both the state-of-the-art learning-based method Unicorn [33] and MPEG standard G-PCC-GesTM v3.0 [6], in terms of rate-distortion trade-off.
@@ 2.1 Traditional Point Cloud Video Compression
Figure 1: The overall architecture of U-Motion for attribute.
@@ 2.2 Learning-Based Point Cloud Video Compression
Early research on learning-based point cloud compression focuses on static point clouds [29, 30, 10, 35, 11]. Wang et al. [29] first propose to use sparse convolution [5] to compress voxelized point clouds, and propose a multiscale architecture, as well as a local auto-regressive architecture SOPA [30] for lossless geometry compression. To further exploit the temporal redundancy in point cloud video, some methods adopt temporal convolution for inter prediction. Anique et al. [2] design a predictor for geometry compression, which compensates and compresses the current frame’s latent representation based on its reference neighbors. Wang et al. introduce Unicorn [32, 33], a unified architecture that integrates geometry & attributes, static & dynamic compression, using context compensation and multiscale motion compensation. The motion compensation is obtained through sparse convolution on the nearby voxels in the previous frame, known as target convolution. However, both Unicorn and Anique et al. lack explicit motion estimation, leading to sub-optimal temporal prediction. On the contrary, Fan et al. propose D-DPCC [9] with an end-to-end learning-based motion estimation and motion compensation (ME/MC) module. Xia et al. [34] further propose a two-layer ME/MC architecture for coarse-to-fine motion compensation. Jiang et al. [17] introduces multiscale motion estimation, where finer-scale’s motion estimation uses coarser scale’s estimated motion as base motion. However, the above ME/MC-based methods are developed only for point cloud video geometry. In addition, their lack of a hierarchical inter- and intra-prediction structure leads to inferior coding efficiency compared to Unicorn-Geometry.
2.3 Learning-based Video Compression
Prior to learning-based point cloud compression, there have been significant advances in learned video compression [23, 15, 16, 19, 27, 20, 21, 22]. DVC [23] and FVC [15] are the pioneering AI video codecs that perform ME/MC on pixel domain and feature domain, respectively. These works follow a residual-coding framework which encodes the difference between the warped decoded reference frame and the current frame. Li et al. propose DCVC [19], which extends the residual-coding framework to context-coding, using motion-compensated reference frame features as temporal context for coding the current frame features. DCVC-HEM [20] further proposes a parallel checkerboard entropy model supporting variable bit-rate. DCVC-DC [21] and DCVC-FM [22] further incorporate hierarchical quality structure and inter feature modulation to alleviate the error accumulation in the temporal context. However, due to the irregular point geometry and data diversity of point cloud sequences, learning-based video compression methods cannot be directly integrated to point cloud video compression.
@@ 3.1 Overview of Attribute Compression
As shown in Figure 1, U-Motion is a hierarchical architecture consisting of $K+1$ ($K=4$ during our experiments) U-Inter and Spatial Predictive Coding layers, where the $l$-th layer corresponds to the $l$-th level of the point cloud when represented by an octree, with the highest layer $H$ corresponding to the original voxelized point cloud. We denote the point cloud frame (including both geometry coordinate and color attribute for each point at each octree level) at time $t$ and layer $l$ as $p_{l}^{t}$, and its reference (i.e. the decoded previous point cloud frame) as $\hat{p}_{l}^{t-1}$. $p_{l}^{t}$ is expressed as a sparse tensor $\left[G(p_{l}^{t}),A(p_{l}^{t})\right]$, where $G(p_{l}^{t})$ denotes an $N\times 3$ geometry coordinate matrix and $A(p_{l}^{t})$ denotes the corresponding $N_{l}\times 3$ attribute matrix, with $N_{l}$ denoting the total number of points at layer $l$. The encoder fuses $p_{l}^{t}$ and the upper layer’s latent $f_{l+1}^{t}$ (after down-sample) to generate the latent representation $f_{l}^{t}$. The same encoder is also used to generate the reference frame’s latent $\hat{f}_{l}^{t-1}$.
The U-Inter module takes the latents $f_{l}^{t}$ and $\hat{f}_{l}^{t-1}$, as well as the upper-layer motion feature $e_{l+1}^{t}$ (through a top-down connection) and lower-layer decoded motion features $\hat{e}_{l-1}^{t}$ and decoded motion $m_{l-1}^{t}$ (through a bottom-up connection), and outputs the motion feature ${e}_{l}^{t}$ at layer $l$. Furthermore, it compresses ${e}_{l}^{t}$ to generate the motion bit stream $r_{e,l}^{t}$ as well as decoded motion feature $\hat{e}_{l}^{t}$. Thereafter, it decodes the motion field $m_{l}^{t}$ from $\hat{e}_{l}^{t}$ and generates the temporal context $\check{f}_{l}^{t}$, by warping $\hat{f}_{l}^{t-1}$ using ${m}_{l}^{t}$.
The Spatial-Temporal Predictive Coding module encodes and decodes current frame’s latent $f_{l}^{t}$ into $\hat{f}_{l}^{t}$ based on the temporal context $\check{f}_{l}^{t}$ and spatial context $\hat{f}_{l-1}^{t}$. Finally, the decoder module takes $\hat{f}_{l}^{t}$ and the lower layer’s decoded points $\hat{p}_{l-1}^{t}$ to recover the color attributes in $\hat{p}_{l}^{t}$.
3.2 Hierarchical Attribute Encoding and Decoding
@@
The encoder in Fig. 1 maps the current layer $p_{l}^{t}$ and the upper-layer latent $f_{l+1}^{t}$ into the latent representation for the current layer $f_{l}^{t}$. The same encoder is used to map $\hat{p}_{l}^{t-1}$ and $\hat{f}_{l+1}^{t-1}$ into $\hat{f}_{l}^{t-1}$. In contrast to Unicorn, which blocks the latents from flowing between layers, our encoder uses both $p_{l}^{t}$ and $f_{l+1}^{t}$ to generate $f_{l}^{t}$, so that the coarser layer’s latent encapsulates important information of the finer layers. We use sparse convolution for network construction, as with all other modules of the entire network (See supplementary).
Figure 2: The network architecture for U-Inter module.
Attribute Decoder.
We assume the geometry has been decoded losslessly and is known. The decoder module recovers the colors in the $l$-th scale $\hat{p}_{l}^{t}$ based on the reconstructed latent $\hat{f}_{l}^{t}$ and coarser scale’s reconstruction $\hat{p}_{l-1}^{t}$. Specifically, the network estimates a residual attribute $\Delta p_{l}^{t}$ at every layer based on the concatenation of upsampled $\hat{p}_{l-1}^{t}$ and $\hat{f}_{l}^{t}$. The reconstruction is the sum of $\Delta p_{l}^{t}$ and the upsampled $\hat{p}_{l-1}^{t}$. The upsample is realized by trans-pooling.
3.3 U-Inter
@@
As shown in Figure 2, our U-Inter Prediction follows a U-Net structure, consisting of bottom-up (coarse-to-fine) motion estimation, top-down motion feature propagation, bottom-up predictive coding of motion, and finally multi-scale motion compensation (warping based on the decoded motions at different scales).
Figure 3: The network architecture for Spatial-Temporal Predictive Coding. For attribute compression, (b) is used for reconstruction; for geometry compression, (a) is used for reconstruction.
Specifically, we adopt the coarse-to-fine motion estimation widely used in video/point cloud compression [34, 17]. For layer $l$, the network first warps reference $\hat{f}_{l}^{t-1}$ based on the coarse encoder motion $\tilde{m}_{l-1}^{t}$. The convolution block performs spatial convolution on the concatenated feature channels from both warped $\hat{f}_{l}^{t-1}$ and ${f}_{l}^{t}$ (using sparse convolution), and outputs encoder motion feature $\tilde{e}_{l}^{t}$, which encodes the residual motion at layer $l$. This block also decodes $\tilde{e}_{l}^{t}$ into $\tilde{m}_{l}^{t}$ for the motion estimation at the upper scale.
Instead of compressing $\tilde{e}_{l}^{t}$ directly, we propose a motion propagation block that fuses $\tilde{e}_{l}^{t}$ with $\tilde{e}_{l+1}^{t}$ (through a top-down connection), and generates fused motion feature $e_{l}^{t}$, which contains motion information of layers $l:H$. This block enables 1) propagation of high-frequency details to the coarser scale, important for correct motion estimation at each scale; 2) distribution of finer-scale motion across multiple scales, to facilitate more efficient and flexible compression of motion information. In later ablation studies, we will show that when this propagation block is removed, the network often yields degraded motions in part due to the rate constraint. Note that this top-down motion propagation introduces redundancies among $e_{l}^{t}$ across layers. Therefore, the bottom-up motion prediction module contextually detaches the redundancy in $e_{l}^{t}$ that is predictable from the decoded coarser-scale $\hat{e}_{l-1}^{t}$, when compressing $e_{l}^{t}$, and contextually restores $\hat{e}_{l-1}^{t}$ into $\hat{e}_{l}^{t}$ on the decoder side.
Multiscale Motion Decoding and Compensation.
Although [17] introduces coarse-to-fine motion estimation, it only applies motion compensation at the finest scale to code the latent features using temporal prediction at the finest scale. We propose progressive coarse-to-fine motion compensation with multiples scales of compensation and predictive coding. The motion decoding module decodes a residual motion $\Delta m_{l}^{t}$ from motion feature $\hat{e}_{l}^{t}$, which is then added to the upsampled $\hat{m}_{l-1}^{t}$ to generate the motion $\hat{m}_{l}^{t}$.
Group Motion
As shown in [21] for motion estimation in 2D video, different feature channels may experience different motions. Borrowing this idea, we estimate multiple motion groups at each scale. The motion decoding module outputs a motion $\hat{m}_{l}^{t}\in\mathbb{R}^{N_{l}^{t}\times G_{l}\times 3}$, where $N_{l}^{t}$ is the number of points in frame $t$ layer $l$, $G_{l}$ is the number of motion groups of layer $l$. Assuming the latent dimension is $C_{l}$, we divide the $C_{l}$ feature channels evenly into $G_{l}$ groups, and decode from the motion features $\hat{e}_{l}^{t}$ a residual motion flow for each group, $\Delta m_{l,g}^{t}$. In general, we find that increasing the number of groups brings gain, despite the potential motion bit rate increase brought by coding more motion groups. However, increasing the group size results in a quadratic increase in complexity for motion compensation. Consequently, using large motion groups at fine scales leads to a sharp rise in computational complexity. Therefore, we employ motion group merging, utilizing large group sizes at coarse scales and progressively merging and reducing the number of motion groups at finer scales.
To derive the decoded motion at scale $l$ for group $g$, we need to generate the upsampled motion from scale $l-1$, which generally has more groups. We average the upsampled motions for the corresponding groups at scale $l-1$, and add to the the decoded residual motion. That is:
$m^{t}_{l,g}=\frac{1}{G^{\prime}}\sum_{i=G^{\prime}g}^{G^{\prime}(g+1)-1}\uparrow m_{l-1,i}^{t}+\Delta m_{l,g}^{t},$ (1)
where $G^{\prime}=\lceil\frac{G_{l-1}}{G_{l}}\rceil$, $\uparrow$ denotes the upsample operator
For each point with coordinate $s$ in $f_{l}^{t}$, its warped coordinate for channel $c$, $\check{s}_{c}=s+m_{l,s,g}^{t}$, may not have a correspondence in $\hat{f}_{l}^{t-1}$, where $g$ denotes the group number corresponding to channel $c$, $m_{l,s,g}^{t}\in\mathbb{R}^{3}$ denotes the motion at time $t$, layer $l$, coordinate $s$ and group $g$. Therefore, we adopt the 3D Adaptive Weighted Interpolation (3DAWI) algorithm [9] to generate the inter context $\check{f}_{l}^{t-1}$,
$\check{f}_{l,s_{c}}^{t}=\frac{\sum_{v\in\vartheta(\check{s}_{c})}d_{v,\check{s}_{c}}^{-1}\cdot\hat{f}_{l,v}^{t-1}}{\max\left\{\sum_{v\in\vartheta(\check{s}_{c})}d_{v,\check{s}_{c}}^{-1},\alpha\right\}},$ (2)
@@
where $\vartheta(\check{s}_{c})$ denotes the K-nearest neighbors of $\check{s}_{c}$ in $\hat{f}_{l}^{t-1}$, $d_{v,\check{s}_{c}}$ is the distance between $\check{s}_{c}$ and $v$, $\alpha$ is a penalty parameter for points too far from $\check{s}_{c}$.
3.4 Spatial-Temporal Predictive Coding